# Few-shot Species Range Estimation

### v1

@@ Abstract
Knowing where a particular species can or cannot be found on Earth is crucial for ecological research and conservation efforts. By mapping the spatial ranges of all species, we would obtain deeper insights into how global biodiversity is affected by climate change and habitat loss. However, accurate range estimates are only available for a relatively small proportion of all known species. For the majority of the remaining species, we often only have a small number of records denoting the spatial locations where they have previously been observed. We outline a new approach for few-shot species range estimation to address the challenge of accurately estimating the range of a species from limited data. During inference, our model takes a set of spatial locations as input, along with optional metadata such as text or an image, and outputs a species encoding that can be used to predict the range of a previously unseen species in feed-forward manner. We validate our method on two challenging benchmarks, where we obtain state-of-the-art range estimation performance, in a fraction of the compute time, compared to recent alternative approaches.
@@ 1 Introduction
Understanding the spatial distribution of plant and animal species is essential to mitigate the ongoing decline in global biodiversity (Jetz et al., 2019). Monitoring these distributions over time allows us to quantify the effects of climate change, habitat loss, and conservation interventions (Mantyka-pringle et al., 2012). Estimating a species’ spatial distribution typically starts with collecting a set of observations that indicate the locations where the species has been confirmed to be present, or absent. Traditionally, this data is used to train models that can then generate detailed predictions over a spatial region of interest (Elith et al., 2006; Beery et al., 2021). When sufficient data is available, these models enable practitioners to estimate important quantities such as the spatial range (i.e., where a species can be found) or abundance (i.e., the total number of individuals) of a species, in addition to quantifying how these quantities are changing.
@@
Despite the availability of well-established modeling techniques, our current understanding of species’ distributions is extremely limited due to little or no observational data being available for most species. For example, iNaturalist, one of the largest citizen science platform documenting global biodiversity, has collected over 130 million research quality observations for approximately 373,000 species globally (iNaturalist, 2025). However, the data is severely long-tailed, i.e., a small percentage of common species account for the majority of the observations, while many species have very few observations. In fact, over half of the 373,000 species cataloged by iNaturalist have been observed fewer than ten times. This data limitation is amplified by the fact that the vast majority of the several million species that are thought to exist have not yet even been documented by science (Mora et al., 2011). Identifying locations where under-observed species can be found is a time consuming and laborious process, often requiring long expeditions to remote locations to search for species that are hard to find. Consequently, there is a pressing need for computational methods that can reliably estimate the spatial distributions of species using only a small number of observations.
@@
Knowing the range of one species can help predict the range of another due to shared ecological, environmental, and geographic contexts. Recent advances in range estimation have leveraged this idea by training joint models using millions of observations across tens of thousands of species (Cole et al., 2023). However, these models still rely on relatively large numbers of training observations for individual species, which limits their applicability to species with limited observations. In this work, we introduce FS-SINR, a novel Transformer-based model that overcomes this limitation and offers two key advantages over previous approaches. First, we obtain superior performance in the few-shot regime — a scenario that represents the reality for the majority of species, yet remains underexplored in prior work. Second, we make accurate predictions for species not present in the training set without any additional training, which can enable interactive exploration and modeling. At inference time, we only require a set of observed locations for the unseen species to generate reliable range estimates. Furthermore, we show we can flexibly incorporate additional non-geographic context information (e.g., a text summary of the species’ habitat or range preferences or an image of the species) to further improve prediction quality. Figure 1 illustrates how FS-SINR can be used at inference time.
@@
In summary, we make the following core contributions: (i) We introduce FS-SINR, a new approach for few-shot species range estimation. FS-SINR has novel capabilities, including the ability to predict the spatial range of a previously unseen species at inference time without requiring any retraining. (ii) We demonstrate, across the challenging IUCN and S&T benchmark datasets, that FS-SINR achieves state-of-the-art performance in the few-shot setting. (iii) We provide detailed ablation studies and visualizations to highlight the benefits of integrating observational data with textual and visual context, as well as to compare our approach with alternative methods.
@@ 2 Related Work
Species Distribution Modeling. Estimating the spatial distribution of a species is a widely explored topic both in statistical ecology and machine learning (Beery et al., 2021). The goal is to develop models that can predict the distribution of species over space, and possibly time, given sparse observation data. Different machine learning approaches using traditional techniques such as decision trees have been extensively explored (Phillips et al., 2004; Elith et al., 2006). More recently, alternative deep learning-based methods have been introduced (Botella et al., 2018; Mac Aodha et al., 2019; Cole et al., 2023; Kellenberger et al., 2024). One of the strengths of these deep methods is that they can jointly represent thousands of different species inside of the same model and have been shown to improve as more training data is added, even when the data is from different species as in SINR (Cole et al., 2023).
@@
There has also been work investigating different approaches for addressing some of the challenges associated with training and evaluating these models. Examples include attempts to addresses imbalances across species in the training observation data (Zbinden et al., 2024b), methods for sampling pseudo-absence data (Zbinden et al., 2024a), biases in the training locations (Chen & Gomes, 2019), representing location information (Rußwurm et al., 2024), discretizing continuous model predictions (Dorm et al., 2024), active learning approaches (Lange et al., 2023), using additional metadata such as images (Teng et al., 2023; Dollinger et al., 2024; Picek et al., 2024) or text (Sastry et al., 2023, 2025; Hamilton et al., 2024), and designing new evaluation datasets to benchmark performance (Cole et al., 2023; Picek et al., 2024). In our work, we investigate the under-explored few-shot setting, where only limited observations (e.g., fewer than ten) are available for each species at training time.
@@
Few-shot Species Range Estimation. There are several aspects of the species range estimation task in the low data regime that makes it different from other few-shot problems more commonly explored in the literature (Parnami & Lee, 2022; Wang et al., 2020). For one, the input domain is fixed (i.e., all locations on earth), each location can support more than one species (i.e., multi-label instead of multi-class), the label space is much larger (i.e., tens of thousands of species as opposed to hundreds of classes in image classification), and only partial supervision is available (e.g., presence-only data, with no confirmed absences).
@@
Lange et al. (2023) introduced an active learning-based approach for species range estimation which makes predictions based on linear combinations of learned species embeddings and showed its effectiveness in the few-shot regime. LE-SINR (Hamilton et al., 2024) showed that internet sourced free-form text descriptions of species’ ranges can be used to train models for zero-shot range estimation. They applied their approach to the few-shot setting, but it requires retraining a classifier for each new species observation added. In our evaluation, we demonstrate that our FS-SINR approach, which can incorporate additional metadata at training time and does not require any retraining during inference, outperforms existing methods.
@@ 3 Methods
Figure 2: FS-SINR overview. Here we depict our few-shot species range estimation model. The input consists of an arbitrary number of context locations $\mathcal{C}^{t}$, that are each independently tokenized using a location encoder $f_{\bm{\theta}}()$, and optional auxiliary context information like text or an image. A class token (CLS) is also appended to the input. All input tokens are processed by a Transformer $m_{\bm{\psi}}()$. Given the set of input context locations, we estimate the probability that a species is present at a query location $\mathbf{x}$ by multiplying the location encoder’s embedding of $\mathbf{x}$ with the projected embedding of the CLS token which is output from the species decoder.
@@ 3.1 Species Range Estimation
We start by describing the SINR approach from Cole et al. (2023). Let $\bm{x}=(\text{lat},\text{lon})\in\mathcal{X}$ be a location of interest sampled from a spatial domain $\mathcal{X}$ (e.g., the surface of the earth). Our goal is train a model $g():\mathcal{X}\to[0,1]^{s}$ to predict the probability of $s$ different species of interest occurring at $\bm{x}$. We let $\hat{\bm{y}}=g(\bm{x})$, where $\hat{y}_{j}\in[0,1]$ (the $j^{th}$ entry of $\hat{\bm{y}}$) represents the probability that species $j$ occurs at location $\bm{x}$.
We can decompose the model as $g()=h_{\bm{\phi}}()\circ f_{\bm{\theta}}()$, where $f_{\bm{\theta}}():\mathcal{X}\rightarrow\mathbb{R}^{d}$ is a location encoder with parameters $\bm{\theta}$ and $h_{\bm{\phi}}():\mathbb{R}^{d}\rightarrow[0,1]^{s}$ is a multi-label classifier with parameters $\bm{\phi}$. The location encoder $f_{\bm{\theta}}()$ maps a location $\bm{x}$ to a $d$-dimensional latent embedding $f_{\bm{\theta}}(\bm{x})$. The multi-label classifier $h()$ is implemented as a per-species linear projection followed by an element-wise sigmoid non-linearity, meaning that $\hat{\bm{y}}=\sigma(f_{\bm{\theta}}(\bm{x})\bm{W})$, where $\bm{W}\in\mathbb{R}^{d\times s}$ (i.e., $h_{\bm{\phi}}()=\bm{\phi}=\bm{W}$) and $\sigma()$ is the sigmoid function. Thus, each column vector $\bm{w}_{j}$ of $\bm{W}$ can be viewed as a species embedding, which we can combine with a location embedding $f_{\bm{\theta}}(\bm{x})$ via an inner product to compute the probability that the species $j$ is present at $\bm{x}$. Importantly, the location embedding is shared across all species. Once trained, it is then possible to generate a prediction for a given species for all locations of interest (e.g., the entire surface of the earth) by evaluating the model at all locations (i.e., $\bm{x}\in\mathcal{X}$).
One of the main challenges associated with training models for species range estimation is that there is a dramatic asymmetry in the available training data. Specifically, it is much easier to collect presence observations (i.e., confirmed sightings of a species) than absence observations (i.e., confirmation that a species is not present at a specific location). As a result, many methods have been developed to train models using presence-only data. In the presence-only setting, we have access to training pairs $(\bm{x},z)$, where $\bm{x}$ is a geographic location and $z\in\{1,\dots,s\}$ is an integer indicating which species was observed there. To overcome the lack of confirmed absence data, one common approach is to generate pseudo-absences by sampling random locations on the surface of the earth (Phillips et al., 2009). Give these pseudo-absences, the parameters of $g()$ can be trained in an end-to-end manner using variants of the cross entropy loss. Specifically, we use the full assume negative loss from Cole et al. (2023) to train the SINR baseline:
@@
where $z$ is the index of the species present for a given training instance, $\hat{y}_{j}$ is the predicted probability of the presence of species $j$, $\hat{y}_{j}^{\prime}$ is the model prediction for a randomly sampled pseudo-absence location, and hyperparameter $\lambda$ balances the presence and pseudo-absence loss terms.
@@ 3.2 Few-shot Range Estimation
We address this challenge by proposing a new approach for few-shot species range estimation called FS-SINR. Our model can predict the probability of presence for a previously unobserved species directly at inference time given only the set of confirmed presence locations available, without any retraining or parameter updates. At inference time we assume we have access to a set of context locations $\mathcal{C}^{t}=\{\bm{c}_{1},\dots,\bm{c}_{k}\}$, which represent a set of $k$ locations where the species $j$ has been confirmed to be present. Each entry is this set denotes a geographic location, i.e., $\bm{c}=(\text{lat},\text{lon})$. Like SINR, our model is also conditioned on a location $\bm{x}$ of interest (i.e., the ‘query’ location), but uses the context locations to inform the prediction for the query location. Note, context locations can come from a species not previously observed by FS-SINR during training.
We represent our FS-SINR model as $g(\bm{x})=m_{\bm{\psi}}(f_{\bm{\theta}}(\bm{x}),\mathcal{C}^{t})$. Unlike in SINR, where the classifier head $h_{\bm{\phi}}()$ is a simple multi-label classifier and sigmoid non-linearity, in our case, the ‘head’ of the model $m_{\bm{\psi}}()$ is a Transformer-based encoder (Vaswani et al., 2017). FS-SINR takes an unordered set of context locations $\mathcal{C}^{t}$ as input, where each location is encoded into an embedding vector (i.e., a token) via a SINR-style multi-layer perceptron location encoder – see Figure 2 for an illustration. Importantly, our model is invariant to the number and ordering of the context locations as we do not append any positional embeddings. This flexibility ensures that it can process a variable number of context locations at inference. We also append an additional register token (REG) as in Darcet et al. (2024) to provide the model with an additional token to ‘store’ information. Given that the input sequence is unordered and may or may not include additional context information, we add learned ‘embedding type’ vectors to each token such that the Transformer knows if a given input token is a location, register, text, image, etc.
@@
FS-SINR uses a similar training loss to $\mathcal{L}_{\mathrm{AN-full}}$. However as it has no equivalent to $h_{\bm{\phi}}()$ we cannot easily include all species in the loss, and instead consider only those within the same batch of training examples $S^{b}$. We obtain a predicted species embedding vector for a given species during the forward pass which can be used to estimate the probabilities of presence for that species for all locations sampled in the batch. We denote this new loss as $\mathcal{L}_{\mathrm{AN-full-b}}$, which indicates that we are considering only those elements contained within the current batch $b$:

### v2

# Feedforward Few-shot Species Range Estimation

@@ Abstract
Knowing where a particular species can or cannot be found on Earth is crucial for ecological research and conservation efforts. By mapping the spatial ranges of all species, we would obtain deeper insights into how global biodiversity is affected by climate change and habitat loss. However, accurate range estimates are only available for a relatively small proportion of all known species. For the majority of the remaining species, we typically only have a small number of records denoting the spatial locations where they have previously been observed. We outline a new approach for few-shot species range estimation to address the challenge of accurately estimating the range of a species from limited data. During inference, our model takes a set of spatial locations as input, along with optional metadata such as text or an image, and outputs a species encoding that can be used to predict the range of a previously unseen species in a feedforward manner. We evaluate our approach on two challenging benchmarks, where we obtain state-of-the-art range estimation performance, in a fraction of the compute time, compared to recent alternative approaches.
@@ 1 Introduction
Understanding the spatial distribution of plant and animal species is essential to mitigate the ongoing decline in global biodiversity (Jetz et al., 2019). Monitoring these distributions over time allows us to quantify the impacts of climate change, habitat loss, and conservation interventions (Mantyka-pringle et al., 2012). Estimating a species’ spatial distribution typically starts with collecting a set of observations that denote the locations where the species has been confirmed to be present or absent. Traditionally, this data is used to train models that can then generate detailed predictions over a spatial region of interest (Elith et al., 2006; Beery et al., 2021). When sufficient data is available, these models enable practitioners to estimate important quantities such as the spatial range (i.e., where a species can be found) or abundance (i.e., the total number of individuals) of a species, in addition to quantifying how these quantities are changing over time.
@@
Despite the availability of well-established modeling techniques, our current understanding of species’ distributions is extremely limited as little or no observational data is available for most species. For example, iNaturalist, one of the largest citizen science platform documenting global biodiversity, has collected over 130 million research quality observations for approximately 373,000 species globally (iNaturalist, 2025). However, the data is severely long-tailed, i.e., a small percentage of common species account for the majority of the observations, while many species have very few observations. In fact, over half of the 373,000 species cataloged by iNaturalist have been observed fewer than ten times to date. This data limitation is amplified by the fact that the vast majority of the several million species that are thought to exist have not yet even been documented by science (Mora et al., 2011). Identifying locations where under-observed species can be found is a time-consuming and laborious process, often requiring long expeditions to remote locations to search for species that are hard to find. Consequently, there is a pressing need for computational methods that can reliably estimate the spatial distributions of species using only a small number of observations.
@@
Knowing the range of one species can help predict the range of another due to shared ecological, environmental, and geographic contexts. Recent advances in range estimation, such as Spatial Implicit Neural Representations (SINR), have leveraged this idea by training on millions of observations, across tens of thousands of species, inside one model (Cole et al., 2023). However, these models still rely on relatively large numbers of training observations for individual species, which limits their applicability to species with limited observations. In this work, we introduce Few-shot Spatial Implicit Neural Representations (FS-SINR), a novel Transformer-based model that overcomes this limitation and offers two key advantages over previous approaches. First, we obtain improved performance in the few-shot regime, a scenario that represents the reality for the majority of species, yet remains underexplored in prior work. Second, we make accurate predictions for species not present in the training set without any additional training, which can enable interactive exploration and modeling. At inference time, we only require a set of observed locations for the unseen species to generate reliable range estimates. Furthermore, we show we can flexibly incorporate additional non-geographic context information (e.g., a text summary of the species’ habitat or range preferences or an image of the species) to further improve prediction quality. Figure 1 illustrates how FS-SINR can be used at inference time.
@@
In summary, we make the following contributions: (i) We introduce FS-SINR, a new approach for few-shot species range estimation. FS-SINR has novel capabilities, including the ability to predict the spatial range of a previously unseen species at inference time without requiring any retraining. (ii) We demonstrate that FS-SINR achieves state-of-the-art performance in the few-shot setting on the challenging IUCN and S&T benchmark datasets. (iii) We provide detailed ablation studies and visualizations to highlight the benefits of integrating observational data with textual and visual context, as well as to compare our approach with alternative methods.
@@ 2 Related Work
Species Distribution Modeling. Estimating the spatial distribution of a species is a widely explored topic in both statistical ecology and machine learning (Beery et al., 2021). The goal is to develop models that can predict the distribution of species over space, and possibly time, given sparse observation data. Different machine learning approaches, initially using traditional techniques, such as decision trees among others have been extensively explored, e.g., (Phillips et al., 2004; Elith et al., 2006). More recently, deep learning-based methods have been introduced (Botella et al., 2018; Mac Aodha et al., 2019; Cole et al., 2023; Kellenberger et al., 2024). One of the strengths of these deep methods is that they can jointly represent thousands of different species within the same model and have been shown to improve as more training data is added, even when the data is from different species (Cole et al., 2023).
@@
There has also been work investigating different approaches to address some of the challenges associated with training and evaluating these models. Examples include attempts to address imbalances across species in the training observation data (Zbinden et al., 2024b), sampling pseudo-absence data (Zbinden et al., 2024a), biases in training locations (Chen & Gomes, 2019), representing location information (Rußwurm et al., 2024), discretizing continuous model predictions (Dorm et al., 2024), active learning approaches (Lange et al., 2023), using additional metadata such as images (Teng et al., 2023; Dollinger et al., 2024; Picek et al., 2024) or text (Sastry et al., 2023, 2025; Hamilton et al., 2024), and designing new evaluation datasets to benchmark performance (Cole et al., 2023; Picek et al., 2024). In our work, we investigate the underexplored few-shot setting, where only limited observations (e.g., fewer than ten) are available for each species at training time.
@@
Few-shot Species Range Estimation. There are several aspects of the species range estimation task in the low-data regime that make it different from other few-shot problems more commonly explored in the literature (Parnami & Lee, 2022; Wang et al., 2020). For example, the input domain is fixed (i.e., all locations on earth), each location can support more than one species (i.e., multi-label instead of multi-class), the label space is much larger (i.e., tens of thousands of species as opposed to hundreds of classes in image classification), and only partial supervision is available (e.g., presence-only data, with no confirmed absences).
Figure 2: FS-SINR overview. Here we depict our few-shot species range estimation model. The input consists of an arbitrary number of context locations $\mathcal{C}^{t}$ for target species $t$ that are each independently tokenized using a location encoder $f_{\bm{\theta}}()$, and optional auxiliary context information like text or an image. A class token (CLS) is also appended to the input. All input tokens are processed by a Transformer $m_{\bm{\psi}}()$. Given the set of input context locations, we estimate the probability that a species is present at a query location $\mathbf{x}$ by multiplying the location encoder’s embedding of $\mathbf{x}$ with the projected embedding of the CLS token which is output from the species decoder.
@@
Lange et al. (2023) introduced an active learning-based approach for species range estimation which makes predictions based on linear combinations of learned species embeddings and showed its effectiveness in the few-shot regime. LE-SINR (Hamilton et al., 2024) showed that internet sourced free-form text descriptions of species’ ranges can be used when training models for zero-shot range estimation. They applied their approach to the few-shot setting, but it requires retraining a classifier for each new species observation added. In our evaluation, we demonstrate that our FS-SINR approach, which can incorporate additional metadata at training time and does not require retraining during inference, outperforms existing methods.
@@ 3.1 Species Range Estimation
We start by describing the SINR approach from Cole et al. (2023). Let $\bm{x}=(\text{lat},\text{lon})\in\mathcal{X}$ be a location of interest sampled from a spatial domain $\mathcal{X}$ (e.g., a location on earth). Our goal is to train a model $g():\mathcal{X}\to[0,1]^{s}$ to predict the probabilities of $s$ different species of interest occurring at $\bm{x}$. We let $\hat{\bm{y}}=g(\bm{x})$, where $\hat{y}_{j}\in[0,1]$ (i.e., the $j^{th}$ entry of $\hat{\bm{y}}$) represents the probability that species $j$ occurs at location $\bm{x}$.
We can decompose the model as $g()=h_{\bm{\phi}}()\circ f_{\bm{\theta}}()$, where $f_{\bm{\theta}}():\mathcal{X}\rightarrow\mathbb{R}^{d}$ is a location encoder with parameters $\bm{\theta}$ and $h_{\bm{\phi}}():\mathbb{R}^{d}\rightarrow[0,1]^{s}$ is a multi-label classifier with parameters $\bm{\phi}$. The location encoder $f_{\bm{\theta}}()$ maps a location $\bm{x}$ to a $d$-dimensional latent embedding $f_{\bm{\theta}}(\bm{x})$. The multi-label classifier $h()$ is implemented as a per-species linear projection followed by an element-wise sigmoid non-linearity, meaning that $\hat{\bm{y}}=\sigma(f_{\bm{\theta}}(\bm{x})\bm{W})$, where $\bm{W}\in\mathbb{R}^{d\times s}$ (i.e., $h_{\bm{\phi}}()=\bm{\phi}=\bm{W}$) and $\sigma()$ is the sigmoid function. Thus, each column vector $\bm{w}_{j}$ of $\bm{W}$ can be viewed as a species embedding, which we can combine with a location embedding $f_{\bm{\theta}}(\bm{x})$ via an inner product to compute the probability that the species $j$ is present at $\bm{x}$. Importantly, the location embedding is shared across all species. Once trained, it is possible to generate a prediction for a given species for all locations of interest by evaluating the model for all locations (i.e., $\bm{x}\in\mathcal{X}$).
One of the main challenges associated with training models for species range estimation is that there is a dramatic asymmetry in the available training data. Specifically, it is much easier to collect presence observations (i.e., confirmed sightings of a species) compared to absence observations (i.e., confirmation that a species is not present at a specific location). As a result, many methods have been developed to train models using presence-only data. In the presence-only setting, we have access to training pairs $(\bm{x},z)$, where $\bm{x}$ is a geographic location, and $z\in\{1,\dots,s\}$ is an integer indicating which species was observed there. To overcome the lack of confirmed absence data, one common approach is to generate pseudo-absences by sampling random locations on the surface of the earth (Phillips et al., 2009). Given these pseudo-absences, the parameters of $g()$ can be trained in an end-to-end manner using variants of the cross-entropy loss. Specifically, we use full assume negative loss from Cole et al. (2023) to train the SINR baseline:
@@
where $z$ is the index of the species present for a given training instance, $\hat{y}_{j}$ is the predicted probability of the presence of species $j$, $\hat{y}_{j}^{\prime}$ is the model prediction for a randomly sampled pseudo-absence location, and the hyperparameter $\lambda$ balances the presence and pseudo-absence loss terms.
@@ 3.2 Few-shot Range Estimation
We address this challenge by proposing a new approach for few-shot species range estimation called FS-SINR. Our model can predict the probability of presence for a previously unobserved species directly at inference time given only the set of confirmed presence locations available, without any retraining or parameter updates. At inference time, we assume we have access to a set of context locations $\mathcal{C}^{t}=\{\bm{c}_{1},\dots,\bm{c}_{k}\}$, which represent a set of $k$ locations where the species $j$ has been confirmed to be present. Each entry in this set denotes a geographic location, i.e., $\bm{c}=(\text{lat},\text{lon})$. Like SINR, our model is also conditioned on a location $\bm{x}$ of interest (i.e., the ‘query’ location), but uses the context locations to inform the prediction for the query location. Note, the context locations can come from a species not previously observed during training.
We represent our model as $g(\bm{x})=m_{\bm{\psi}}(f_{\bm{\theta}}(\bm{x}),\mathcal{C}^{t})$. Unlike in SINR, where the classifier head $h_{\bm{\phi}}()$ is a simple multi-label classifier and sigmoid non-linearity, in our case, the ‘head’ of the model $m_{\bm{\psi}}()$ is a Transformer-based encoder (Vaswani et al., 2017). FS-SINR takes an unordered set of context locations $\mathcal{C}^{t}$ as input, where each location is encoded into an embedding vector (i.e., a token) via a SINR-style multi-layer perceptron location encoder – see Figure 2 for an illustration. Importantly, our model can accept a variable number of context locations and is invariant to their ordering as we do not append any positional embeddings. This flexibility ensures that it can process a variable number of context locations during inference. We also append an additional register token (REG) as in Darcet et al. (2024) to provide the model with an additional token to ‘store’ information. Given that the input sequence is unordered and may or may not include additional context information, we add learned ‘embedding type’ vectors to each token such that the Transformer knows if a given input token is a location, register, text, image, etc.
@@
FS-SINR uses a similar training loss to $\mathcal{L}_{\mathrm{AN-full}}$. However, since it has no equivalent to $h_{\bm{\phi}}()$ we cannot easily include all species in the loss, and instead consider only those within the same batch of training examples of size $s^{b}$. We obtain a predicted species embedding vector for a given species during the forward pass which can be used to estimate the probabilities of presence of that species for all locations sampled in the batch. We denote this new loss as $\mathcal{L}_{\mathrm{AN-full-b}}$, which indicates that we are considering only those elements contained within the current batch $b$: